OpenStudy (anonymous):

A square has a perimeter of 36 inches. What is the area? (1 points) 81 in2 72 in2 9 in2 6 in2

OpenStudy (minecrafter64):

just do 36 times 36

OpenStudy (anonymous):

Square has 4 equal sides so to find length of 1 side is Perimeter/ 4.

OpenStudy (anonymous):

and area of square is side squared there is answer

OpenStudy (anonymous):

\[P=4a\] \[A=a^2\] Using these equations where P is perimeter, A is area, a is side length, calculate your side length and hence find area

OpenStudy (anonymous):

9 squared?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

p = 4s 36 = 4s s = 36/4 s = 9

OpenStudy (anonymous):

Area of a square = s^2

OpenStudy (anonymous):

so, now whats 9^2 in other words what is 9 times 9?

OpenStudy (anonymous):

its 81